# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

### v2

@@ Abstract
The deployment of extremely large antenna arrays (ELAAs) and operation at higher frequency bands in wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) systems introduce significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties. Combined with dual-wideband impacts, these effects fundamentally reshape the sparsity patterns of wideband XL-MIMO channels in the angular-delay domain, making existing sparsity-based channel estimation methods inadequate. To address these challenges, this paper revisits the channel estimation problem for wideband XL-MIMO systems, considering dual-wideband effects, spherical wavefront, and SnS properties. By leveraging the spatial-chirp property of near-field array responses, we quantitatively characterize the sparsity patterns of wideband XL-MIMO channels in the angular-delay domain, revealing global block sparsity and local common-delay sparsity. Building on this structured sparsity, we formulate the wideband XL-MIMO channel estimation problem as a multiple measurement vector (MMV)-based Bayesian inference task and propose a novel column-wise hierarchical prior model to effectively capture the sparsity characteristics. To enable efficient channel reconstruction, we develop an MMV-based variational message passing (MMV-VMP) algorithm, tailored to the complex factor graph induced by the hierarchical prior. Simulation results validate the proposed algorithm, demonstrating its convergence and superior performance compared to existing methods, thus establishing its effectiveness in addressing the challenges of wideband XL-MIMO channel estimation under complex near-field conditions.
@@ I Introduction
Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2]. By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[3]. It also delivers substantial beamforming gains, effectively mitigating the severe path loss characteristic of millimeter-wave (mmWave) frequencies. Additionally, the availability of GHz-wide bandwidth in the mmWave spectrum alleviates spectrum congestion, reinforcing the potential of mmWave XL-MIMO for next-generation wireless networks [4]. However, realizing these performance benefits critically depends on obtaining accurate channel state information (CSI), underscoring the importance of precise channel knowledge.
@@ I-A Related Works
Due to the limited scattering resulting from the highly directional propagation behavior of mmWave, mmWave channels typically exhibit significant sparsity. Consequently, most existing channel estimation methods are designed to exploit this inherent sparsity structure to simplify and improve the estimation process. For instance, leveraging the sparsity in the polar domain, various channel estimation and beam training schemes have been proposed to address the spherical effects in XL-MIMO channels [14, 15, 16]. Considering the inherent spatial-domain sparsity induced by SnS properties, several Bayesian inference-based methods have been developed [17, 18] to mitigate the SnS effects. Additionally, exploiting the joint sparsity of XL-MIMO channels in both the spatial and angular domains, methods such as [8, 19] have proposed two-stage and joint VR detection and channel estimation algorithms, considering both spherical wavefront and SnS properties. However, despite these advancements, these methods predominantly focus on narrowband systems without the considerations of dual-wideband effects.
@@
More recently, methods addressing both the spherical wavefront effect and dual-wideband effects have been proposed. For example, [25] introduced a bilinear pattern detection (BPD)-based approach to recover wideband XL-MIMO channels, while [26] developed a hybrid message passing algorithm based on constrained Bethe free energy minimization. However, despite these advancements, none of these approaches explicitly incorporate SnS properties, highlighting a significant gap in wideband XL-MIMO channel estimation.
@@ I-B Motivations and Contributions
  • To accurately characterize the SnS dual-wideband XL-MIMO channels, we derive a novel spatial-frequency domain model based on the measurement findings reported in [11], which comprehensively accounts for the spherical wavefront effect, SnS characteristics, and dual-wideband effects. By leveraging the spatial-chirp property of the array response, we quantitatively investigate the impact of the spherical wavefront effect, SnS properties, and dual-wideband effects on the angular and delay spreads. Furthermore, we reveal the sparsity patterns of SnS dual-wideband XL-MIMO channels in the angular-delay domain, incorporating both global block sparsity and local common-delay sparsity.
  • Exploiting the sparsity of the angular-delay domain, the SnS dual-wideband XL-MIMO channel estimation is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem. Recognizing the critical role of an accurate prior model in Bayesian inference, we propose a novel column-wise hierarchical prior model, where Markov random field (MRF) and common-delay sparsity mechanism is introduced to capture both the global block sparsity and the local common-delay sparsity, ensuring a more accurate and structured representation of the channel characteristics.
  • To enable efficient Bayesian inference, we propose a MMV-based variational message passing (MMV-VMP) algorithm tailored for the column-wise hierarchical prior model. Unlike the variance states propagation algorithm in [27], which relies on first-moment matching methods under the mean-field approximation to update prior model parameters, the MMV-VMP algorithm employs structured variational inference combined with belief propagation. This approach effectively handles the higher-order dependencies and loops in the complex factor graph, ensuring precise and reliable message updates.
  • Compared to existing methods, the proposed MMV-VMP algorithm achieves superior performance by jointly exploiting the global and common-delay sparsity of SnS dual-wideband channels in the angular-delay domain. Additionally, as the algorithm avoids matrix inversion operations, it maintains moderate computational complexity, making it well-suited for practical deployment in XL-MIMO systems.
@@ I-C Organization and Notations
The rest of the paper is organized as follows: Section II introduces the mmWave XL-MIMO system model. Section III develops a spatial-frequency channel model, incorporating spherical wavefront effects, SnS properties, and dual-wideband effects, and analyzes the angular-delay characteristics of SnS dual-wideband XL-MIMO channels. Section IV formulates the channel estimation problem as a sparse signal recovery task and proposes a column-wise hierarchical sparse prior model. Section V presents the MMV-VMP algorithm. Simulation results and conclusions are provided in Sections VI and VII, respectively.
@@ II System Model
Consider a mmWave XL-MIMO system, as shown in Fig. 1, where a base station (BS) utilizes an ELAA comprising $N$ antennas arranged in a uniform linear configuration to serve a single-antenna user. To reduce hardware cost and energy consumption, the BS adopts a hybrid analog-digital architecture, where $N_{\mathrm{RF}}$ radio frequency (RF) chains are connected to the ELAA through phase shifters. The spacing between adjacent antennas is $d=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$, with $c$ and $f_{c}$ indicating the speed of light and the central carrier frequency, respectively. In addition, to address frequency-selective fading, OFDM is utilized. Without loss of generality, assume that $K$ subcarriers are uniformly selected from the available subcarriers to carry pilot symbols.
In this paper, we focus on the uplink channel estimation, where the BS combines the pilot signal using all RF chains associated with different beams. Let the total number of received beams be denoted by $M$, with $M$ being an integer multiple of $N_{\mathrm{RF}}$. Consequently, the BS requires $P=M/N_{\mathrm{RF}}$ channel uses to cycle through all beams for a given pilot symbol. Denote $\iota_{k,p}$ as the pilot symbol of the $k$-th subcarrier in the $p$-th channel use. Then, the received signal for the $k$-th subcarrier in the $p$-th channel use can be expressed as
@@ III-A Channel Modeling for SnS XL-MIMO Systems
Assume that there are $L$ propagation paths between the BS and user. Denote =$\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l>1$ indicates reflection or diffraction path, as illustrated in Fig. 1. Denote $\tau_{l,n}$ as the time delay of the $l$-th path for the $n$-th antenna. Then the baseband channel impulse response of the $n$-th received antenna is given by
$h_{n}(t)=\sum_{l=1}^{L}\tilde{\alpha}_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l,n}}\delta(t-\tau_{l,n}),$ (4)
@@
where $(a)$ is obtained by using the Fresnel approximation, i.e., $\sqrt{1+x}\approx 1+x/2-x^{2}/8$. In this Taylor expansion, the current point is set at $n=0$, and terms of third order and higher are neglected. Utilizing (7), (6) can be rewritten as $\tau_{l,n}=\tau_{l}+n{\psi_{l}}/{f_{c}}-n^{2}{\varphi_{l}}/{f_{c}}$, where $\psi_{l}\triangleq d\cos\vartheta_{l}/\lambda_{c}$ and $\varphi_{l}\triangleq{d^{2}\sin^{2}\vartheta_{l}}/{2r_{l}\lambda_{c}}$. According to (6) and (7), the continuous time channel impulse response in (4) can be further expressed as
$h_{n}(t)=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\delta\left(t-\tau_{l,n}\right),$ (8)
where $\alpha_{l}\triangleq\tilde{\alpha}_{l}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l}}$ is the equivalent complex path gain. Applying the continuous time Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (10), as shown in the top of this page.
$h_{n}(f)=\int_{-\infty}^{+\infty}h_{n}(t)\mathrm{e}^{-\mathrm{j}2\pi ft}\mathrm{d}t=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}\mathrm{e}^{-\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}.$ (10)
@@
where $a(f,\tau_{l})=\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}$; $\boldsymbol{\kappa}_{l}=[\kappa_{l,1},\kappa_{l,2},\cdots,\kappa_{l,N}]^{\mathrm{T}}$ denotes the visibility indicator vector with the $n$-th entry being $\kappa_{l,n}$; $\mathbf{b}(\psi_{l},\varphi_{l})$ and $\mathbf{a}(f,\tau_{l})\}$ denote the array response vector and the frequency response vector, respectively, with each of their $n$-th entries given by
$[\mathbf{b}(\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{-\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)},$ (12)
$[\boldsymbol{\theta}(f,\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{-\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}.$ (13)
Assuming the total bandwidth of the OFDM system is denoted by $f_{s}$, the center frequency for each subcarrier is given by $f_{k}=f_{c}+{kf_{s}}/{K}$ with $k=0,1,\cdots,K$. The corresponding channel vector for each subcarrier is then represented as $\mathbf{h}(f_{k})$, as illustrated in Eq. (11) with $f$ replaced by $f_{k}$. Thus, the overall spatial-frequency channel matrix $\mathbf{H}$ can be reformulated as
@@
where $\mathbf{a}(\tau_{l})=[a(f_{1},\tau_{l}),\cdots,a(f_{K},\tau_{l})]^{\mathrm{T}}$, and $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})=[\boldsymbol{\theta}(f_{1},\psi_{l},\varphi_{l}),\cdots,\boldsymbol{\theta}(f_{K},\psi_{l},\varphi_{l})]$.
@@ III-B Angular-Delay Representation of SnS XL-MIMO Channel
$\mathbf{H}\approx\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}.$ (15)
@@
where $\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{N\times N}$ and $\mathbf{F}_{\mathrm{D}}\in\mathbb{C}^{K\times K}$ denote the normalized $N$- and $K$-dimensions DFT matrices. $\mathbf{X}\in\mathbb{C}^{N\times K}$ denotes the angular-delay channel.
Lemma 1. The angular-delay channel $\mathbf{X}=\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\mathbf{H}\mathbf{F}_{\mathrm{D}}^{*}$ is a sparse matrix containing only $L$ non-zero blocks, with the non-zeros elements staying inside a square region, i.e.,

### v3

@@ Abstract
The deployment of extremely large antenna arrays (ELAAs) in wideband extremely large-scale multiple-input multiple-output (XL-MIMO) systems introduces significant near-field effects, such as spherical wavefront propagation and spatially non-stationary (SnS) properties. When combined with the dual-wideband effects inherent to wideband systems, these phenomena fundamentally reshape the channel’s sparsity patterns in the angular-delay domain, rendering existing sparsity-based channel estimation methods inadequate. To address these challenges, this paper reconsiders the channel estimation problem for wideband XL-MIMO systems. By leveraging the spatial-chirp property of array responses, we quantitatively characterize the angular-delay domain sparsity of wideband XL-MIMO channels, revealing both global block sparsity and local common-delay sparsity. To effectively capture this structured sparsity, we propose a novel column-wise hierarchical prior model that incorporates precision parameter sharing mechanism and Markov random field (MRF) structure. Based on this prior model, the channel estimation task is formulated as a multiple measurement vector (MMV)-based Bayesian inference problem. Tailored to the complex factor graph induced by this hierarchical prior, we develop an MMV-based hybrid message passing (MMV-HMP) algorithm. The algorithm performs message updates along the factor graph edges by selectively applying either the sum-product algorithm (SPA) or variational message passing (VMP) rules, depending on the factor-node structure and message tractability. Simulation results validate the effectiveness of the proposed column-wise hierarchical prior model through comprehensive ablation studies and demonstrate that the MMV-HMP algorithm consistently outperforms existing baselines, achieving robust estimation performance with moderate computational complexity.
@@ I Introduction
Wideband extremely large-scale multiple-input-multiple- output (XL-MIMO) has emerged as a pivotal technology to meet the capacity demands of future 6G communications [1, 2, 3]. By exploiting the spatial degrees of freedom (DoFs) provided by extremely large antenna arrays (ELAAs), XL-MIMO significantly enhances both spectral and energy efficiencies[4, 5]. It also delivers substantial beamforming gains, effectively mitigating the severe path loss characteristic of millimeter-wave (mmWave) frequencies. Additionally, the availability of GHz-wide bandwidth in the mmWave spectrum alleviates spectrum congestion, reinforcing the potential of mmWave XL-MIMO for next-generation wireless networks [6, 7]. However, realizing these performance benefits critically depends on obtaining accurate channel state information (CSI), underscoring the importance of precise channel knowledge.
@@ I-A Related Works
Due to the limited scattering resulting from the highly directional propagation behavior of mmWave, mmWave channels typically exhibit significant sparsity. Consequently, most existing channel estimation methods are designed to exploit this inherent sparsity structure to simplify and improve the estimation process. For instance, leveraging the sparsity in the polar domain, various channel estimation and beam training schemes have been proposed to address the spherical effects in XL-MIMO channels [17, 18, 19]. Considering the inherent spatial-domain sparsity induced by SnS properties, several Bayesian inference-based methods have been developed [20, 21] to mitigate the SnS effects. Additionally, exploiting the joint sparsity of XL-MIMO channels in both the spatial and angular domains, methods such as [11, 22] have proposed two-stage and joint visibility region (VR) detection and channel estimation algorithms, considering both spherical wavefront and SnS properties. However, despite these advancements, these methods predominantly focus on narrowband systems without the considerations of dual-wideband effects.
@@
More recently, methods addressing both the spherical wavefront effect and dual-wideband effects have been proposed. For example, [28] introduced a bilinear pattern detection (BPD)-based approach to recover wideband XL-MIMO channels, while [29] developed a message passing algorithm based on constrained Bethe free energy minimization. However, despite these advancements, none of these approaches explicitly incorporate SnS properties, highlighting a significant gap in wideband XL-MIMO channel estimation.
@@ I-B Motivations and Contributions
  • To accurately characterize the SnS dual-wideband XL-MIMO channels, we derive a spatial-frequency domain model that comprehensively captures the spherical wavefront propagation, SnS characteristics, and dual-wideband effects. By exploiting the spatial-chirp property of array responses, we quantitatively analyze the influence of spherical wavefront propagation, SnS properties, and dual-wideband effects on the angular and delay spreads. Furthermore, we reveal the resulting sparsity patterns in the angular-delay domain, which exhibit both global block sparsity and local common-delay sparsity.
  • Exploiting the sparsity in the angular-delay domain, the SnS dual-wideband XL-MIMO channel estimation problem is formulated as a multiple measurement vector (MMV)-based sparse recovery task. Recognizing the importance of an accurate prior model in channel reconstruction, we introduce a novel column-wise hierarchical prior, which incorporates both a precision parameter sharing mechanism and a Markov random field (MRF) structure. This design enables the model to effectively capture global block sparsity and local common-delay sparsity, thereby providing a more structured and accurate representation of the underlying channel characteristics.
  • To achieve efficient Bayesian inference on the intricate factor graph, resulting from the incorporation of both the precision parameter sharing mechanism and the MRF structure, we propose a MMV-based hybrid message passing (MMV-HMP) algorithm that jointly leverages belief propagation (BP) and variational message passing (VMP) rules. Thanks to this flexible message update strategy, the MMV-HMP algorithm can effectively manage the message updates for latent precision parameters with non-Gaussian distributions and accommodate the loops inherent in the complex factor graph, thereby enabling accurate and reliable inference.
  • Compared to existing approaches, the proposed MMV-HMP algorithm achieves superior performance by jointly capturing the global block sparsity and local common-delay sparsity inherent in SnS dual-wideband XL-MIMO channels. This enables robust channel estimation while maintaining moderate computational complexity. In addition, the effectiveness of the proposed column-wise hierarchical prior model is validated through comprehensive ablation studies.
@@ I-C Organization and Notations
The rest of the paper is organized as follows: Section II introduces the mmWave XL-MIMO system model. Section III develops a spatial-frequency channel model, incorporating spherical wavefront effects, SnS properties, and dual-wideband effects, and analyzes the angular-delay characteristics of SnS dual-wideband XL-MIMO channels. Section IV formulates the channel estimation problem as a sparse signal recovery task and proposes a column-wise hierarchical sparse prior model. Section V presents the MMV-HMP algorithm. Simulation results and conclusions are provided in Sections VI and VII, respectively.
@@ II System Model
Consider a mmWave XL-MIMO OFDM system, as shown in Fig. 1, where a base station (BS) utilizes an ELAA comprising $N$ antennas arranged in a uniform linear configuration. To reduce hardware cost and energy consumption, the BS adopts a hybrid analog-digital architecture, where $N_{\mathrm{RF}}$ radio frequency (RF) chains are connected to the ELAA through phase shifters. The spacing between adjacent antennas is $d=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$, with $c$ and $f_{c}$ indicating the speed of light and the central carrier frequency, respectively. For the uplink channel estimation, we assume the $J$ users transmit mutual orthogonal pilot sequences to the BS [30, 26, 31], e.g., orthogonal time or frequency resources are utilized for different users to transmit pilot sequences. Therefore, channel estimation for each user is independent. Without loss of generality, we consider an arbitrary user and assume $K$ subcarriers are uniformly selected from the available subcarriers to carry pilot symbols.
@@
In the uplink channel estimation phase, the BS combines the pilot signal using all RF chains associated with different beams. Let the total number of received beams be denoted by $M$, with $M$ being an integer multiple of $N_{\mathrm{RF}}$. Consequently, the BS requires $P=M/N_{\mathrm{RF}}$ channel uses to cycle through all beams for a given pilot symbol. Denote $\iota_{k,p}$ as the pilot symbol of the $k$-th subcarrier in the $p$-th channel use. Then, the received signal for the $k$-th subcarrier in the $p$-th channel use can be expressed as
@@ III-A Channel Modeling for SnS XL-MIMO Systems
Assume that there are $L$ propagation paths between the BS and user. Denote =$\mathcal{L}=\left\{1,2,\cdots,L\right\}$ as the set of all propagation paths, where $l=1$ refers to the line-of-sight (LoS) path and $l\geq 2$ indicates reflection or diffraction path, as illustrated in Fig. 1. Denote $\tau_{l,n}$ as the time delay of the $l$-th path for the $n$-th antenna. Then the time-invariant baseband channel impulse response of the $n$-th received antenna can be given by[32]
$h_{n}(\tau)=\sum_{l=1}^{L}\tilde{\alpha}_{l}\kappa_{l,n}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l,n}}\delta(\tau-\tau_{l,n}),$ (4)
@@
where $(a)$ is obtained by using the Fresnel approximation, i.e., $\sqrt{1+x}\approx 1+x/2-x^{2}/8$. In this Taylor expansion, the current point is set at $n=0$, and terms of third order and higher are neglected. Utilizing (7), (6) can be rewritten as $\tau_{l,n}=\tau_{l}-n{\psi_{l}}/{f_{c}}+n^{2}{\varphi_{l}}/{f_{c}}$, where $\psi_{l}\triangleq d\cos\vartheta_{l}/\lambda_{c}$ and $\varphi_{l}\triangleq{d^{2}\sin^{2}\vartheta_{l}}/{2r_{l}\lambda_{c}}$. According to (6) and (7), the channel impulse response in (4) can be further expressed as
$h_{n}(\tau)=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\delta\left(\tau-\tau_{l,n}\right),$ (8)
where $\alpha_{l}\triangleq\tilde{\alpha}_{l}\mathrm{e}^{-\mathrm{j}2\pi f_{c}\tau_{l}}$ is the equivalent complex path gain. Applying the Fourier transform to (8), the spatial-frequency response of the $n$-th antenna is given by (9)[33, 34], as shown in the top of this page.
$h_{n}(f)=\int_{-\infty}^{+\infty}h_{n}(\tau)\mathrm{e}^{-\mathrm{j}2\pi f\tau}\mathrm{d}\tau=\sum_{l=1}^{L}\alpha_{l}\kappa_{l,n}\mathrm{e}^{\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)}\mathrm{e}^{\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}.$ (9)
@@
where $a(f,\tau_{l})=\mathrm{e}^{-\mathrm{j}2\pi f\tau_{l}}$; $\boldsymbol{\kappa}_{l}=[\kappa_{l,1},\kappa_{l,2},\cdots,\kappa_{l,N}]^{\mathrm{T}}$ denotes the visibility indicator vector with the $n$-th entry being $\kappa_{l,n}$; $\mathbf{b}(\psi_{l},\varphi_{l})$ and $\boldsymbol{\theta}(f,\psi_{l},\varphi_{l})$ denote the array response vector and the frequency-dependent phase-shift vector, respectively, with each of their $n$-th entries given by
$[\mathbf{b}(\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{\mathrm{j}2\pi\left(n{\psi_{l}}-n^{2}{\varphi_{l}}\right)},$ (11)
$[\boldsymbol{\theta}(f,\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{\mathrm{j}2\pi f\left(n\frac{\psi_{l}}{f_{c}}-n^{2}\frac{\varphi_{l}}{f_{c}}\right)}.$ (12)
Assuming the total bandwidth of the OFDM system is denoted by $f_{s}$, the center frequency for each pilot subcarrier is given by $f_{k}=f_{c}+{kf_{s}}/{K}$ with $k=0,1,\cdots,K-1$. The corresponding channel vector for each subcarrier is then represented as $\mathbf{h}(f_{k})$, as illustrated in Eq. (10) with $f$ replaced by $f_{k}$. Thus, the overall spatial-frequency channel matrix $\mathbf{H}$ can be reformulated as
@@
where $\mathbf{a}(\tau_{l})=[a(f_{1},\tau_{l}),\cdots,a(f_{K},\tau_{l})]^{\mathrm{T}}$, and $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})=[\boldsymbol{\theta}(f_{0},\psi_{l},\varphi_{l}),\cdots,\boldsymbol{\theta}(f_{K-1},\psi_{l},\varphi_{l})]$.
@@ III-B Angular-Delay Representation of SnS XL-MIMO Channel
$\mathbf{H}=\mathbf{F}_{\mathrm{A}}\mathbf{X}\mathbf{F}_{\mathrm{D}}^{\mathrm{T}},$ (14)
@@
where $\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{N\times N}$ and $\mathbf{F}_{\mathrm{D}}\in\mathbb{C}^{K\times K}$ denote the normalized $N$- and $K$-dimensions discrete Fourier transformation (DFT) matrices. $\mathbf{X}\in\mathbb{C}^{N\times K}$ denotes the angular-delay channel.
From (11) and (12), it can be observed that both the frequency-dependent and frequency-independent array response vectors take the form of spatial chirp signals. Consequently, they can be reformulated as
$[\mathbf{b}(\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{\mathrm{j}2\pi\left({\psi_{l}}-n{\varphi_{l}}\right)n},$ (15)
$[\boldsymbol{\theta}(f_{k},\psi_{l},\varphi_{l})]_{n}$ $=\mathrm{e}^{\mathrm{j}2\pi\left(a_{l,k}n-\frac{b_{l,k}}{2}n\right)n},$ (16)
where $a_{l,k}=\psi_{l}{f_{k}}/{f_{c}}$ and $b_{l,k}=2\varphi_{l}{f_{k}}/{f_{c}}$ denote the initial spatial frequency and the spatial chirp rate, respectively.
By exploiting the chirp-like properties described in (15), the start and end spatial frequencies of the SnS array response vector $\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})$ can be approximated as [35]
$i_{l,s}=\psi_{l}-2\varphi_{l}{n_{l,s}},\quad i_{l,e}=\psi_{l}-2\varphi_{l}{n_{l,e}}.$ (17)
In this manner, the array response vector $\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})$, which captures both the spherical wavefront effect and the SnS characteristics, can be approximated as a superposition of multiple far-field array response components, i.e.,
$\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\int_{i_{l,e}}^{i_{l,s}}c(\epsilon_{l})\mathbf{b}_{\mathrm{far}}(\epsilon_{l})\mathrm{d}\epsilon_{l},$ (18)
where $c(\epsilon_{l})$ denotes the decomposition coefficient, and $[\mathbf{b}_{\mathrm{far}}(\epsilon_{l})]_{n}=\mathrm{e}^{\mathrm{j}2\pi\epsilon_{l}n}$ represents the far-field array response corresponding to spatial frequency $\epsilon_{l}\in[i_{l,e},i_{l,s}]$. Furthermore, the continuous decomposition in (18) can be discretized with a spatial frequency resolution of $2/N$ as
$\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\sum_{p=1}^{P}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p}),$ (19)
where the number of sampling points is given by $P=\lceil({n_{l,s}}-{n_{l,e}})\varphi_{l}N\rceil$, and the discretized spatial frequency is defined as $\epsilon_{l,p}=\min(i_{l,s},i_{l,e})+{2(p-1)}/{N}$ for $p=\{1,2,\dots,P\}$. The coefficient $c_{l,p}$ denotes the corresponding weight associated with the $p$-th far-field array response component.
By utilizing the decomposition in (19), the spatial-frequency channel described in (13) can be expressed as
$\mathbf{H}=\sum_{l=1}^{L}\sum_{p=1}^{P}\alpha_{l}c_{l,p}\mathbf{b}_{\mathrm{far}}(\epsilon_{l,p})\mathbf{a}^{\mathrm{T}}(\tau_{l})\odot\boldsymbol{\Theta}(\psi_{l},\varphi_{l}).$ (20)
Similar to the analysis in (17), the start and end spatial frequencies of the SnS array response vector $\boldsymbol{\theta}(f_{k},\psi_{l},\varphi_{l})$ can be approximated as
$i_{l,s}(k)$ $=(a_{k}-b_{k}n_{s}){\lambda_{\mathrm{c}}}/{d},$ (21)
$i_{l,e}(k)$ $={(a_{k}-b_{k}n_{e})}\lambda_{\mathrm{c}}/{d}.$ (22)
In addition, for notation simplification, we define
$t_{l,n}=\frac{n^{2}{f_{s}\varphi_{l}}-n{f_{s}\psi_{l}}}{Kf_{c}}.$ (23)
Lemma 1. By utilizing equations (20)-(22), it can be shown that the two-dimensional inverse discrete Fourier transform (2D-IDFT) of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ results in a sparse matrix, where the nonzero entries are confined within a square region. Within this region, the dominant components exhibit a clear column-wise clustered sparsity structure. Specifically, as $N,K\rightarrow\infty$, the support of the nonzero region converges to: